### v1

# Robust Tail Index Estimation under Random Censoring via Minimum Density Power Divergence Nour Elhouda Guesmia, Abdelhakim Necir∗, Djamel

Abstract
Based on the minimum density power divergence approach, we propose a robust estimator of the tail index for randomly right-censored data from a Pareto-type distribution. We establish its consistency and asymptotic normality. An extensive simulation study is performed to assess the finite-sample behavior of the estimator in comparison with existing ones. The methodology is further illustrated through an application to a real AIDS survival dataset.
Keywords: Asymptotic normality; Heavy-tails; robust estimation; Random right censoring; Tail index.
AMS 2020 Subject Classification: 62G32; 62G05; 62G20; 62G35.
Corresponding author:
ah.necir@univ-biskra.dz
E-mail address:
nourelhouda.guesmia@univ-biskra.dz (N. Guesmia)
djamel.meraghni@univ-biskra.dz (D. Meraghni)
1. Introduction
Right-censored data pose a common challenge in statistics, as values are only known up to a certain threshold, while those beyond it remain unobserved. This issue frequently arises in fields such as survival analysis, reliability engineering, and insurance. Pareto-type distributions are commonly used to model data exhibiting extreme values. These distributions are characterized by a relatively high probability of producing observations far from the mean, in contrast to the normal distribution. They are crucial in applications like insurance, finance, and environmental studies, where rare events can have a substantial impact. Let be a sample of size from a random variables (rv) and let be another sample from a rv defined on a probability space with continuous cumulative distribution functions (cdf’s) and respectively. Assume that and are independent. Also, we suppose that is right censored by that is each for stage we can only observe the variable and the indicator variable which determines whether or not has been observed. We assume that the tails functions and are regularly varying at infinity (or Pareto-type) of positive tail indices and respectively. On the other terms for any we have
We denote the cdf of by Then, by the independence between and we have which yields that is also regularly varying at infinity, with tail index In the presence of extreme values and right-censored data, several methods have been proposed for estimating the tail index. Numerous techniques have been developed to address the challenges arising from the nature of such data. Many studies have focused on modifying traditional tail index estimation methods, such as Hill’s estimator (Hill, 1975) to make them more suitable for censored data. Einmahl et al. (2008) adapted this estimator to handle the characteristics of right-censored data, and it is given by where denotes the well-known Hill estimator corresponding to the tail index and
stands for an estimator of the proportion of upper non-censored observations
The integer sequence represents the number of top order statistics used in the estimation of the tail index so that and as The sequence of rv’s represents the order statistics pertaining to the sample and denotes the corresponding concomitant values, satisfying for such that On the basis of a Kaplan-Meier integration, Worms and Worms (2014) proposed a consistent estimator to the tail index defined by
where denotes the well-known Kaplan-Meier estimator of cdf (Kaplan and Meier, 1958). The asymptotic normality of is established in Beirlant et al. (2019) by considering Hall’s model (Hall, 1982). The bias reduction in tail index estimation for censored data is addressed by Beirlant et al. (2016), Beirlant et al. (2018) and Beirlant et al. (2019). By using a Nelson-Aalen integration, recently Meraghni et al. (2025) also derived a new estimator for given by
where
denotes the well-known Nelson-Aalen estimator of cdf (Nelson, 1972). Note that
thus the above formula of may be rewritten into
where
The authors showed that the two tail index estimators and exhibit similar performances in terms both of bias and mean squared error. Actually, the studies made on the comparison between the Kaplan–Meier and Nelson–Aalen estimators reached that the latter exhibit an almost similar statistical behavior; see, for instance, Colosimo et al. (2002). On the other hand, reaching to establish the asymptotic properties of the extreme Kaplan–Meier integrals poses somewhat difficulties. To overcome this issue, Meraghni et al. (2025) introduced a Nelson–Aalen tail-product limit process and established its Gaussian approximation in terms of standard Wiener processes. Thus, they established the consistency and asymptotic normality of the proposed estimator by considering the first and second order conditions of regular variation functions, namely the assumptions and respectively. As expected, they also showed that both the asymptotic bias and variance of and are respectively equal.
The most common method for estimating the parameters of an extreme value distribution in extreme value analysis comes from the maximum likelihood estimation method. Although these estimators possess desirable asymptotic properties, they may be sensitive to outlying observations from the assumed extreme value models; see, for instance, Brazauskas and Serfling (2000). Then, robust statistics are a better approach to avoid outliers and therefore deviations from the underlying parametric models. It has been shown that employing robust statistical ideas in extreme value theory improves the quality and precision of estimates (Dell’Aquila and Embrechts, 2006). Juárèz and Schucany (2004) seem to be the first authors to employ the MDPD of Basu et al. (1998) for the robust estimation of the parameters of an extreme value distribution. Since then, this divergence measure has become the most sought-after divergence measure for robust estimation of parameters of extreme value distributions. Kim and Lee (2008), Dierchx et al. (2013), Goegebeur et al. (2014), Dierckx et al. (2021) have made use of the MDPD in estimating the tail index and quantiles from Pareto-type distributions. Recently, Ghosh (2017) proposed a robust MDPD estimator for a real-valued tail index. This estimator is a robust generalization of the estimator proposed by Mattys and Beirlant (2003) and addresses the non-identical distributions of the exponential regression model using the approach in Ghosh and Basu (2013). Also, Dierchx et al. (2013) employs the MDPD concept on an extended Pareto distribution for relative excesses over a high threshold. Minkah el al. (2023A) and Minkah el al. (2023B) proposed a robust estimator for the tail index of a Pareto-type using the MDPD approach on an exponential regression model.
The remainder of this paper is organized as follows. Section 2 provides a brief overview of the MDPD estimation method, originally introduced by Basu et al. (1998). Section 3 discusses the asymptotic properties of the proposed estimator, namely its consistency and asymptotic normality, with detailed proofs deferred to Section 6. The finite-sample performance of the estimator is examined in Section 5 through a simulation study, including comparisons with existing estimators. Section 7 collects several technical lemmas and propositions. Section 8 presents the figures corresponding to the simulation results.
2. Minimum density power divergence
Given two probability densities and Basu et al. (1998) introduced a new measure of divergence between them, known as the density power divergence:
where is a tuning parameter that controls the trade-off between efficiency and robustness. The case corresponding to is obtained as a limiting case of the general expression as yielding the classical Kullback–Leibler divergence, denoted Let us consider a parametric family of densities and suppose that the goal is to estimate the parameter Let denote the cdf corresponding to the density The minimum density power divergence (MDPD) is a functional defined as The term in does not depend on and thus does not affect the minimization of over Therefore, the minimization involved in computing the MDPD functional reduces to minimizing the objective function
Given a random sample from the cdf we estimate by replacing with the empirical distribution For a given tuning parameter the MDPD estimator of is then defined as the minimizer (over the quantity
Equivalently, this can be written in terms of the observed data as:
Thus, the parameter that minimizes satisfies the estimating equation:
The tuning parameter plays a key role by balancing the trade-off between efficiency and robustness in the MDPD estimator. Specifically, when is close to zero, the estimator tends to be highly efficient but less robust to outliers. As increases, robustness improves at the expense of efficiency. Studies have shown that small values of provide strong robustness properties with only a minor loss in asymptotic efficiency compared to the maximum likelihood estimator under correct model specification.
2.1. MDPD estimator of for right-censored data
Let us consider the relative excess rv given with cdf of corresponding density function Recall that then for large Then we may assume that the rv has the density
The density power divergence objective function corresponding to this model is given by:
By letting and replacing with the Nelson–Aalen estimator we obtain the empirical version:
This defines the empirical objective function Thereby, the MDPD estimator of is then obtained by minimizing That is, it solves the equation:
For we have
Hence, we obtain:
Meraghni et al. (2025) stated that where and are the empirical counterparts of the cdf and the sub-distribution respectively, defined by:
See, for example, Shorack and Wellner (1986) page 294. Here, denotes the left limit of the function at By substituting with the above expression and using the ratio from we obtain:
where is the random weight sequence defined in The MDPD tail index estimator under random censoring is defined as the solution to this equation, which has not yet been addressed in the literature. It is worth mentioning that, for the solution of coincides with the estimator introduced earlier.
3. Consistency and asymptotic normality
From now on, we denote the true value of by
Theorem 3.1.
(consistency) Assume that the cdf’s and satisfy the first-order condition and that Let us be a sequence of integers such that and as then, for any with probability tending to there exists a solution to the estimating equation such that as
Since weak approximations of extreme value theory-based statistics are typically achieved within the second-order framework (see de Haan and Stadtmüller, 1996), it is therefore natural to assume that the cdf satisfies the well-known second-order condition of regular variation. Specifically, we assume that for any
Here is the second-order parameter, and is a function tending to not changing sign near infinity, and whose absolute value is regularly varying at infinity with index When the expression is interpreted as In what follows, the functions and respectively denote the quantile and tail quantile functions of a given cdf For convenience, we set and define for
Theorem 3.2.
(asymptotic normality) Assume that cdf’s satisfies the second-order condition and that Let us be a sequence of integers such that and and Then for any
as where and are respectively defined in and
4. Simulation study
In this section, we evaluate the performance and robustness of the proposed estimator by comparing it to the estimator To assess the robustness of , we consider an -contaminated model where the distribution function is defined as a mixture:
with denoting the contamination proportion. We examine scenarios of mixed censoring based on the Burr distribution with parameters i.e.
contaminated by another Burr distribution with parameters i.e.
Censoring is implemented using a Fréchet distribution with parameter defined as
In our simulation, we fix and consider two values for namely and The proportion is set to and with computed using the relation To assess the effect of outliers, we choose when and when The robustness parameter is varied across four levels: (corresponding to and to evaluate increasing degrees of robustness. We also consider three contamination levels: (no contamination), and For each contamination scenario, we generate samples of size from the specified mixture model. The estimators are compared in terms of absolute bias (Bias) and mean squared error (MSE), evaluated as a functions of the number of upper order statistics. Appendix B presents the graphical results, with Figures 8.2, 8.3 and 8.4 corresponding to the case and Figures 8.5, 8.6 and 6 corresponding to the case .
Turning to the graphical results, we first analyze the effects of both the contamination level and the robustness parameter on the estimator When the estimator demonstrates relative stability, even as the contamination level increases. However, for higher values of the proportion increasing the robustness parameter . In contrast, when the influence of the robustness parameter becomes more pronounced, particularly for and In these cases, larger values of result in notable improvements in estimator performance in terms of both bias and mean squared error (MSE). Overall, the simultaneous increase in and tends to deteriorate the performance of compared to Moreover, in most settings, increasing the number of upper order statistics used in the estimation, leads to higher bias and MSE for.
5. Real data application
The Australian AIDS survival dataset, provided by Dr. P. J. Solomon and the Australian National Centre in HIV Epidemiology and Clinical Research, includes data on patients diagnosed with AIDS in Australia prior to July For our analysis, we focus on the male patients within this dataset, which has been extensively studied in the literature, for instance, Ripley and Solomon (1994) and Venables and Ripley (2002) (pages 379-385), Einmahl et al. (2008), Ndao et al. (2014) and Stupfler (2016). In the context of locally robust estimation of Pareto-type tails, Dierckx et al. (2021) highlighted the effectiveness of the MDPD estimator. They also pointed out that their dataset did not exhibit any notable outliers. They also noted that their dataset did not exhibit any notable outliers. We evaluated the robustness of our proposal by introducing artificial contamination into the dataset. Table 5.1 shows the ten largest observed survival times for male patients replaced by ten outliers.
original time | outlier time | original time | outlier time |
Figure 5.1 presents a comparison of the MDPD estimators of obtained from the original and contaminated datasets, illustrating the effect of contamination. The introduction of ten outliers has a pronounced impact on the estimator (corresponding to as evidenced by a noticeable increase in its estimated value. In contrast, our proposed estimator for exhibits remarkable robustness, maintaining stable estimates despite the presence of these outliers.